OpenStudy (anonymous):

Joe rode from his house to a store at an average speed of 30 kilometers per hour. He walked back along the same route at an average speed of 2 kilometers per hour. If his total traveling time was 2 hours, what is the one- way distance from Joe's house to the store in kilometers?

OpenStudy (anonymous):

\[\frac{d}{30}+\frac{d}{2}=2\] that is, time going plus time returning, using time is distance divided by rate. solve for d

OpenStudy (anonymous):

first maybe add the fractons get \[\frac{d+15d}{30}=2\] \[\frac{16d}{30}=2}\] \[16d=60\] \[d=60\div 16=\frac{15}{4}=37.5\]

OpenStudy (anonymous):

oops \[\frac{16d}{30}=2\]

OpenStudy (anonymous):

and also \[\frac{15}{4}=3.75\] not what i wrote